# Multilingual Political Views of Large Language Models: Identification and Steering

### v1

@@ 3.3 Technical Details
Similar to Röttger et al. (2024), who use a reverse-engineered version of the PCT [^5] https://politicalcompass.github.io/, we apply the same scoring algorithm based on weighted response values. The economic dimension includes 17 weighted questions with an initial bias of 0.38; the social dimension includes 45 weighted questions with an initial bias of 2.41. For normalization, scores are divided by 8.0 (economic) and 19.5 (social), yielding coordinates in the range [10-10- 10, +10+10+ 10] on both axes.
@@ 3.4 Identification Results
There are statistically significant ideological differences between languages within each model. Using the Kruskal–Wallis test Kruskal and Wallis (1952), we find that both the social and economic dimensions differ significantly across languages for every model (see Appendix E for full results). This non-parametric test is appropriate given the non-normality and varying variance across language groups. In some models, such as the Aya-Expanse series, cross-paraphrase variance within each language is low–indicating stable ideological outputs–yet the differences between languages remain substantial. Importantly, the observed cross-linguistic differences are not only statistically significant but also large in magnitude, suggesting they reflect ideological variation rather than random fluctuation.
Model Social (Soc.) Economic (Ec.)
Qwen-3-8B 16 13
Qwen-3-14B 49 31
Qwen-3-32B 43 33
LLaMA-3.1-8B 19 2
LLaMA-3.1-70B 31 25
Aya-Expanse-8B 28 36
Aya-Expanse-32B 53 44
Table 2: Number of statistically significant language pairs per model for social and economic dimensions, based on two-sided Mann–Whitney U tests with Bonferroni correction.
To further examine these differences, we perform pairwise comparisons between all language pairs for each model using the two-sided Mann–Whitney U test Mann and Whitney (1947) with Bonferroni correction Armstrong (2014). This test is well-suited for assessing whether two independent samples differ in distribution, without assuming normality. The results confirm that many language pairs differ significantly in both ideological dimensions, confirming that models exhibit language-specific political behavior. Moreover, the number of statistically significant language pairs increases with model size (Table 2), suggesting that larger models develop more finely differentiated ideological representations across languages. Detailed pairwise results are reported in Appendix E.
@@ 4.1 Test-time Intervention
We attempt to shift the ideological leaning of LLaMA-3.1-8B using the test-time intervention method proposed by Li et al. (2024), similar to Kim et al. (2025). This method operates in two stages. In the first stage, we train binary linear probes on top of the output of every attention head to identify those most responsive to a target classification–in our case, distinguishing between liberal and conservative ideologies. For training the probes, we use the JyotiNayak/political_ideologies dataset from Hugging Face Lhoest et al. (2021), containing 1280 samples for each of the two classes.
[…]
@@
This method leverages the localized semantic capabilities of individual attention heads while being minimally invasive to the model’s overall behavior. We tune two hyperparameters experimentally: $K$ (the number of heads to intervene on) and $\alpha$ (the intervention strength). To isolate the effects of intervention, all generation parameters are deterministic, as specified in Appendix A.
@@ 4.2 Intervention Results
Figure 3 illustrates the effect of test-time intervention on English-language prompted LLaMA-3.1-8B. We apply the intervention on the top 256 most responsive attention heads–constituting one-quarter of all available heads–using a steering vector scaled by an intervention strength of $\alpha=25$. The plot shows the average political compass coordinates across all 11 paraphrased prompts, before and after intervention. The intervention reliably shifts model outputs toward the target ideology across all paraphrases, indicating that targeted manipulation of attention head outputs can steer ideological content in a consistent and interpretable manner. Appendix C provides a comparison across two values of $K$ and three values of $\alpha$.
Figure 3: Intervention results for LLaMA-3.1-8B for $K$=256 and $\alpha$=25. Top: direction towards conservative. Bottom: direction towards liberal.
We further test whether steering vectors derived from English texts can generalize to other languages. Specifically, we apply them to model responses in Turkish (tr), Romanian (ro), Slovenian (sl), and French (fr), and find that they are indeed effective in shifting ideological outputs across languages. Detailed results are provided in Appendix C.
@@
$\alpha$=20 0.0 0.6$\pm$0.8 0.5$\pm$0.5 19.8$\pm$5.4 19.5$\pm$5.3 2.2$\pm$1.3 0.18$\pm$0.4 0.6$\pm$0.7 19.6$\pm$9.1 5.2$\pm$2.8
$\alpha$=25 0.0 1.4$\pm$1.2 0.09$\pm$0.3 32.8$\pm$2.9 25.6$\pm$4.8 4.5$\pm$2.9 0.0 1.1$\pm$1.5 17.5$\pm$8.8 14.5$\pm$5.8
$\alpha$=30 0.0 0.7$\pm$0.8 0.0 30.2$\pm$6.4 37.1$\pm$6.5 34.5$\pm$9.6 0.0 11.5$\pm$9.3 10.5$\pm$6.4 10.3$\pm$5.5
Table 3: Average useless response counts by intervention configuration for various languages. $K$ is set to 256 for all $\alpha$ values. Base results are noted under each language. Cl. 0 and 1 refer to steering towards conservative and liberal directions, respectively.
@@ 6 Conclusion
Our study presents several limitations. Certain models (e.g., LLaMA-3.1) refuse to answer some portions of questions despite forced choice constraints, limiting result completeness. We force responses rather than using free-text evaluation, which may not capture natural model behavior. Our intervention experiments use limited hyperparameter exploration and focus only on LLaMA-3.1-8B, restricting generalizability across architectures.
The Political Compass Test, while widely adopted, represents a Western-centric framework that may inadequately capture political orientations across diverse cultural contexts. We evaluate only instruction-tuned models, which undergo extensive alignment that may mask underlying base model biases. The intervention methodology employs a single steering technique with probes trained only on English-language data. Finally, models were trained on data with different temporal cutoffs, making direct comparisons potentially confounded by evolving political discourse.
@@ Appendix C Political Intervention Results
(c) Intervention Strength of 30
(d) Intervention Strength of 30
Figure 5: Political compass intervention results on 256 heads for two different intervention strengths for both directions on the PCT test in English. Interestingly, the plots on the right demonstrate steering towards politically left responses, and the plots on the left–towards politically right responses.
@@
Figure 6: Political compass intervention results on 512 heads for two different intervention strengths for both directions on the PCT test in English. Interestingly, the plots on the right demonstrate steering towards politically left responses, and the plots on the left–towards politically right responses.
@@
Figure 7: Political compass intervention results on 256 heads for two different intervention strengths for both directions on the PCT test in Romanian. Interestingly, the plots on the right demonstrate steering towards politically left responses, and the plots on the left–towards politically right responses.
@@
Figure 8: Political compass intervention results on 256 heads for two different intervention strengths for both directions on the PCT test in Turkish. Interestingly, the plots on the right demonstrate steering towards politically left responses, and the plots on the left–towards politically right responses.
@@
Figure 9: Political compass intervention results on 256 heads for two different intervention strengths for both directions on the PCT test in Slovenian. Interestingly, the plots on the right demonstrate steering towards politically left responses, and the plots on the left–towards politically right responses.
@@
Figure 10: Political compass intervention results on 256 heads for two different intervention strengths for both directions on the PCT test in French. Interestingly, the plots on the right demonstrate steering towards politically left responses, and the plots on the left–towards politically right responses.

### v2

@@ 3.3 Technical Details
Similar to Röttger et al. (2024), who use a reverse-engineered version of the PCT [^5] https://politicalcompass.github.io/, we apply the same scoring algorithm based on weighted response values. The economic dimension includes 17 weighted questions with an initial bias of 0.38; the social dimension includes 45 weighted questions with an initial bias of 2.41. For normalization, scores are divided by 8.0 (economic) and 19.5 (social), yielding coordinates in the range [10-10, +10+10] on both axes.
@@ 3.4 Identification Results
There are statistically significant ideological differences between languages within each model. Using the Kruskal–Wallis test Kruskal and Wallis (1952), we find that both the social and economic dimensions differ significantly across languages for every model (see our GitHub repository for full results). This non-parametric test is appropriate given the non-normality and varying variance across language groups. In some models, such as the Aya-Expanse series, cross-paraphrase variance within each language is low–indicating stable ideological outputs–yet the differences between languages remain substantial. Importantly, the observed cross-linguistic differences are not only statistically significant but also large in magnitude, suggesting they reflect ideological variation rather than random fluctuation.
Model Social Economic
Qwen-3-8B 16 (17.6%) 13 (14.3%)
Qwen-3-14B 49 (53.8%) 31 (34.1%)
Qwen-3-32B 43 (47.3%) 33 (36.3%)
LLaMA-3.1-8B 19 (20.9%) 2 (2.2%)
LLaMA-3.1-70B 31 (34.1%) 25 (27.5%)
Aya-Expanse-8B 28 (30.8%) 36 (39.6%)
Aya-Expanse-32B 53 (58.2%) 44 (48.4%)
Table 2: Number of statistically significant language pairs per model for social and economic dimensions. Percentages (in parentheses) are relative to the total number of possible language pairs ($n=91$) across 14 languages.
To further examine these differences, we perform pairwise comparisons between all language pairs for each model using the two-sided Mann–Whitney U test Mann and Whitney (1947) with Bonferroni correction Armstrong (2014). This test is well-suited for assessing whether two independent samples differ in distribution, without assuming normality. The results confirm that many language pairs differ significantly in both ideological dimensions, confirming that models exhibit language-specific political behavior. Moreover, the number of statistically significant language pairs increases with model size (Table 2), suggesting that larger models develop more finely differentiated ideological representations across languages.
@@ 4.1 Test-time Intervention
We attempt to shift the ideological leaning of LLaMA-3.1-8B using the test-time intervention method proposed by Li et al. (2024), similar to Kim et al. (2025). This method operates in two stages. In the first stage, we train binary linear probes on top of the output of every attention head to identify those most responsive to a target classification–in our case, distinguishing between liberal and conservative ideologies. For training the probes, we use the JyotiNayak/political_ideologies dataset from Hugging Face [^7] https://huggingface.co/datasets/JyotiNayak/political_ideologies Lhoest et al. (2021), which contains 1,280 samples per class (liberal and conservative) generated by GPT-4 AI (2023). Each sample consists of a short paragraph (2–3 sentences) labeled with one of the two ideological categories. The dataset covers a diverse range of topics, including economic, environmental, family and gender issues, geopolitics and foreign policy, political institutions, racial justice and immigration, religion, and social, health, and education policies.
In the second stage, we intervene on the $K$ most responsive attention heads at inference time. For each selected head, we modify its output by adding a scaled steering vector, $\alpha\cdot\vec{v}$, where $\alpha$ controls the strength of the intervention. These steering vectors are constructed using a simple center-of-mass method [^8] The center-of-mass method Li et al. (2024), originally related to whitening and coloring transformations (Ioffe and Szegedy, 2015; Huang and Belongie, 2017) in deep learning, is equivalent to the DiffMean approach that has emerged in the steering literature (Marks and Tegmark, 2023; Wu et al., 2025)., as opposed to directly using probe weights as in Kim et al. (2025): for each attention head, we compute the mean activation vectors (centers-of-mass) for the liberal and conservative classes, normalize them, and define the direction vector $\vec{v}$ as the normalized difference between these two means. Specifically, for a given attention head $(l,h)$, the steering direction is:
@@
This method leverages the localized semantic capabilities of individual attention heads while remaining minimally invasive to the model’s overall behavior. [^9] Following Li et al. (2024); Panickssery et al. (2023); Arditi et al. (2024), interventions on model activations of this type have been shown not to impair general capabilities. Based on this prior evidence, we do not evaluate this aspect explicitly, as our primary goal is to assess whether political bias can be instilled; evaluation of general capabilities is left for future work. We tune two hyperparameters experimentally: $K$ (the number of heads to intervene on) and $\alpha$ (the intervention strength). To isolate the effects of the intervention, all generation parameters are kept deterministic, as described in Appendix A.
@@ 4.2 Intervention Results
Figure 3 illustrates the effect of test-time intervention on English-language prompted LLaMA-3.1-8B. We apply the intervention on the top 512 most responsive attention heads–constituting a half of all available heads–using a steering vector scaled by an intervention strength of $\alpha=20$. The plot shows the average political compass coordinates across all 11 paraphrased prompts, before and after intervention. The intervention reliably shifts model outputs toward the target ideology across all paraphrases, indicating that targeted manipulation of attention head outputs can steer ideological content in a consistent and interpretable manner. Appendix C provides a comparison across two values of $K$ and three values of $\alpha$.
Figure 3: Intervention results for LLaMA-3.1-8B for KK=512 and $\alpha$=20. Top: direction towards liberal. Bottom: direction towards conservative.
@@
[…]
Table 3: Average irrelevant response counts by intervention configuration for various languages. $K$ is set to 256 for all $\alpha$ values. Base results are noted under each language. Cl. 0 and 1 refer to steering towards liberal and conservative directions, respectively.
Motivated by the findings of Wendler et al. (2024), who show that multilingual prompts are internally routed through English representations within LLMs, we further test whether steering vectors derived from English texts can generalize cross-linguistically. Specifically, we apply these vectors to model responses in Turkish (tr), Romanian (ro), Slovenian (sl), and French (fr), and find that they are indeed effective in shifting ideological outputs across languages to a certain degree. Detailed results are provided in Appendix C.
@@ 6 Conclusion
Our study presents several limitations. Certain models (e.g., LLaMA-3.1) refuse to answer some portions of questions despite forced-choice constraints, limiting result completeness. We force responses rather than using free-text evaluation, which may not capture natural model behavior. Our intervention experiments use limited hyperparameter exploration and focus only on LLaMA-3.1-8B, restricting generalizability across architectures. The Political Compass Test, while widely adopted, represents a Western-centric framework that may inadequately capture political orientations across diverse cultural contexts. We evaluate only instruction-tuned models, which undergo extensive alignment that may mask underlying base model biases. The intervention methodology employs a single steering technique with probes trained only on English-language data. Additionally, while we intervene on attention heads in this study, interventions could also be applied to other model components, such as MLP outputs or post-residual activations, which may be more effective and warrant further investigation. Finally, models were trained on data with different temporal cutoffs, making direct comparisons potentially confounded by evolving political discourse.
@@ Appendix C Political Intervention Results
(c) Intervention Strength of 25
(d) Intervention Strength of 25
(e) Intervention Strength of 30
(f) Intervention Strength of 30
Figure 5: Political compass intervention results on 256 heads for two different intervention strengths for both directions on the PCT test in English. The plots on the right demonstrate steering towards politically right responses, and the plots on the left–towards politically left responses.
@@
Figure 6: Political compass intervention results on 512 heads for two different intervention strengths for both directions on the PCT test in English. The plots on the right demonstrate steering towards politically right responses, and the plots on the left–towards politically left responses.
@@
Figure 7: Political compass intervention results on 256 heads for two different intervention strengths for both directions on the PCT test in Romanian. The plots on the right demonstrate steering towards politically right responses, and the plots on the left–towards politically left responses.
@@
Figure 8: Political compass intervention results on 256 heads for two different intervention strengths for both directions on the PCT test in Turkish. The plots on the right demonstrate steering towards politically right responses, and the plots on the left–towards politically left responses.
@@
Figure 9: Political compass intervention results on 256 heads for two different intervention strengths for both directions on the PCT test in Slovenian. The plots on the right demonstrate steering towards politically right responses, and the plots on the left–towards politically left responses.
@@
Figure 10: Political compass intervention results on 256 heads for two different intervention strengths for both directions on the PCT test in French. The plots on the right demonstrate steering towards politically right responses, and the plots on the left–towards politically left responses.